OpenStudy (anonymous):

Factor the trinomial, 2x^2 - 2x - 12

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

are you drawing?

OpenStudy (raden):

u may can make like this 2x^2-2x-12=2(x^2-x-6) so, to get factor of x^2-x-6 u must think of 2 number when we multiply both get -6 and when to add we get -1

OpenStudy (raden):

sorry, my conection is low :(

OpenStudy (anonymous):

thas ok

OpenStudy (raden):

did u get it

OpenStudy (anonymous):

x-6, x-1

OpenStudy (raden):

nopes... 2x^2-2x-12=2(x^2-x-6)=2(x...?)(x....?)

OpenStudy (raden):

find factor of -6 ?